### v2

# Janis Zenkner

Abstract
LambdaBeam is a state-of-the-art execution-guided algorithm for program synthesis that incorporates higher-order functions, lambda functions, and iterative loops into the Domain-Specific Language (DSL). LambdaBeam generates every program from the start. Yet, many program blocks or subprograms occur frequently in a given domain, e.g., loops to traverse a list. Thus, repeating programs can be used to enhance the synthesis algorithm. However, LambdaBeam fails to leverage this potential. For this purpose, we introduce AbstractBeam: A novel program synthesis framework that employs Library Learning to identify such program repetitions, integrates them into the DSL, and thus utilizes their potential to boost LambdaBeam’s synthesis algorithm. Our experimental evaluations demonstrate that AbstractBeam significantly improves LambdaBeam’s performance in the LambdaBeam integer list manipulation domain. Additionally, AbstractBeam’s program generation is more efficient compared to LambdaBeam’s synthesis. Finally, our findings indicate that Library Learning is effective in domains not specifically crafted to highlight its benefits.
1 Introduction
Effectively enabling machines to write programs based on a set of specifications provided by a user is the goal in the field of program synthesis [16]. Many techniques can be employed to automatically generate programs, such as (Graph) Neural Networks [1], Sequence-to-Sequence [2], or Large Language Models [7]. In addition to foundational end-to-end program synthesis techniques, employing Domain-Specific Languages (DSLs) and Programming-By-Example (PBE) can refine and target the automated generation of programs. PBE takes a unique approach by allowing program specifications through Input-Output (I/O) examples, from which a synthesis tool generates the corresponding code [29]. DSLs are tailor-made languages designed to address particular problems in specific domains, offering concise syntax and operations optimized for specific tasks [38].
A primitive approach to solve a PBE task is to generate and execute all possible combinations of operations in a DSL. However, the amount of combinations grows exponentially with the program length. This combinatorial space, i.e., the search space for a solution, can be visualized as a decision tree. As the programs get longer, the depth of the tree increases. Similarly, for growing DSLs the breadth of the tree increases. Consequently, the trial-and-error approach becomes infeasible. To address this, many DSL-based program synthesis approaches apply combinatorial searches guided by Artificial Neural Networks to traverse the search space more efficiently [3, 30, 4, 32]. Yet, the search space blow-up that is caused by increased search depths and breaths remains the Achilles heel of DSL-based program synthesis.
Additionally, programming is an inherently complex cognitive task, combining logic, creativity, and systematic problem-solving [26]. Human programmers often develop code using a bottom-up approach [12]. This involves crafting small, self-contained functions that address specific aspects of a larger problem. These functions are then combined to construct the complete software solution. An additional strategy to minimize costs and complexity is the use of pre-built software libraries. For example, in the domain of Machine Learning, practitioners employ a neural model architecture from the PyTorch library, leveraging work already done to streamline their own development process [36].
Inspired by human software development, approaches like DreamCoder [15] incorporate Library Learning into a top-down synthesis. Library Learning refers to the identification and extraction of frequently occurring program repetitions or subprograms to create reusable components for more efficient code generation. DreamCoder integrates this procedure into a wake-sleep cycle: During the wake phase, a simple Bayesian top-down algorithm generates programs and during sleeping repeating subprograms are identified. These code repetitions or abstractions are added to the DSL, constituting an automatic mechanism that enhances preexisting software libraries. As abstractions compromise multiple operations and constants, the length of synthesized programs and thus the search depth is reduced. LambdaBeam [34] compromises only a synthesis algorithm that focuses on narrowing the search breadth. For this purpose, it employs an execution-guided bottom-up search. During bottom-up synthesis, subprograms are executed and the subsequent search is biased on these intermediate program states. As programs can be generated step-by-step the guidance is more fine-grained. Additionally, LambdaBeam incorporates higher-order functions, enhancing its capability to handle complex synthesis tasks.
Even though LambdaBeam’s program synthesis is effective in terms of programs evaluated, it does not use Library Learning. Consequently, LambdaBeam is less competent in terms of runtime. In other words, LambdaBeam only needs to generate a few programs to find a solution yet the generation of these few programs takes comparably long. DreamCoder however, with its focus on Library Learning, utilizes abstractions but lacks the potential of fine-grained contextual information that is present during synthesis. Moreover, the domains DreamCoder was evaluated on contain multiple task clusters that differ only in one parameter (e.g., add-k with k=1, add-k with k=2, …). Thus, abstractions can be easily found and applied to other tasks which makes this synthetic domain perfect for Library Learning but bypasses the scalability and generalization.
This situation presents an opportunity for a more unified approach that can harness the strengths of execution-guided bottom-up synthesis and Library Learning. We postulate that enhancing LambdaBeam with Library Learning leads to a better performance. Additionally, we reason that abstractions tackle the inefficiency in terms of time of LambdaBeam’s search policy by reducing the program length. Consequently, the search depth is reduced which in turn causes the performance boost. Furthermore, we assume that abstractions can be found in domains that were not artificially designed to contain sufficient program repetitions. Finally, we postulate that abstractions do not generalize beyond the domain and are only beneficial if they reflect the domain.
Driven by these hypotheses, we propose a new framework — AbstractBeam — that utilizes the strengths of LambdaBeam’s search policy and a Library Learning framework inspired by DreamCoder. Thus, AbstractBeam can define, use, and reason about operations and abstractions that potentially contain lambda functions, iterative loops, and higher-order functions. We showcase the effectiveness of AbstractBeam in the LambdaBeam list manipulation domain. Moreover, we show the effectiveness of Library Learning in a domain that was not specifically designed to show the potential of program repetitions which gives more insights into the scalability of AbstractBeam. Our AbstractBeam code is in the supplementary materials [^1] The code will be made publicly available after publication..
2 Background
2.1 Programming By Example
Rather than describing requirements in natural language, PBE tasks describe a program in the form of I/O examples. The goal is to find a program $\mathcal{P}\>\epsilon\>\mathcal{L}$ that maps all inputs =$\mathcal{I}=\{I_{1},...,I_{N}\}$ to their respective outputs $\mathcal{O}=\{O_{1},...,O_{N}\}$. Here, $\mathcal{L}$ refers to the space of all possible programs that is spanned by a DSL containing constants, input variables, and operations. Operations can be arbitrarily nested and be applied to constants and input variables to solve a task. Exemplary tasks can be found in Appendix A.
2.2 $\mathbf{\lambda}$-Calculus
The lambda calculus [10] is a formal system in mathematical logic that can simulate any Turing Machine. Terms in the lambda calculus are defined as variables, function applications, and lambda abstractions. These lambda abstractions create new functions that include a set of variables specific to their scope and the body of the function itself. We permit lambda abstractions to introduce multiple variables simultaneously, unlike traditional lambda functions which ’curry’ them. Moreover, we use manually designed, domain-specific primitives, such as Add, Sort, or Map, to extend the lambda calculus. Finally, we use simply-typed lambda calculus [28] aiding in data type management and function validation.
2.3 LambdaBeam
Figure 1 illustrates the overall process of synthesizing programs using LambdaBeam [34]. In the initial step of program generation, the set of explored programs - termed ’values’ - includes only input variables and constants. Using a bottom-up approach, LambdaBeam selects arguments for operations based on their arity, i.e., the number of arguments a function requires. To make informed decisions about which arguments to choose for a function $f$, LambdaBeam utilizes a neural module that assesses the search context. This context comprises the I/O examples for the task, the specific operation $f$, and the previously explored values. The neural module then outputs a probability distribution over these values, guiding the selection of arguments $a_{i}$ for the operation $f$. During training, beam search is used to select arguments according to this distribution. When selected arguments $a_{i}$ are lambda functions LambdaBeam employs the same strategy to fill variables. Because beam search operates deterministically, using it to obtain argument lists as done in training can cause the search to stall. This occurs if all newly generated values are identical to previously encountered values and the set of values remains unchanged. To address this issue, LambdaBeam uses two strategies. First, the UniqueRandomizer [31] mitigates always choosing the same value by methodically selecting unique samples from the model and recording these in a specialized data structure, which prevents the reselection of previously chosen values. Secondly, the search process incorporates periodic restarts at defined intervals instead of a single continuous execution.
Figure 1: LambdaBeam synthesis overview [34]. The synthesis model utilizes contextual information from the task specifications in form of the I/O examples, the current operation, and so far explored values to condition the selection of arguments on.
This method is adopted to counteract the difficulties the model faces in correcting its course from ineffective search paths, as periodic restarts help to alleviate this challenge. However, this approach involves a trade-off where previously explored values are discarded with each restart. After selecting all necessary arguments, the function $f(a_{0},...,a_{arity})$ is executed to generate new outputs, which are then added to the set of explored values. Each new expression’s weight, i.e., its length, is also recorded. This iterative process continues until the algorithm synthesizes a program that meets the given specifications or until the search times out.
Figure 2: LambdaBeam model architecture [34]. The I/O Module processes I/O examples of tasks into a vector $e_{IO}$ using an MLP. The Value Model embeds the set of explored values and their weights employing an MLP. The Search-Context Summary Module combines the value embeddings and I/O vector into a vector representation $e_{c}$. Finally, based on an $LSTM_{op}$ architecture, the Argument Selector Module predicts the distribution used to determine the argument list for a given operation.
The architecture of LambdaBeam’s neural model is structured around four components: the I/O Module, the Value Model, the Search-Context Summary Module, and the Argument Selector Module. Figure 2 shows the individual modules and their interaction. Further details on the learning algorithm of LambdaBeam can be found in Appendix B.
2.4 Stitch
Stitch [5] implements a method for the automated synthesis of abstractions from a set of programs, i.e., a program corpus. Starting with an initial DSL, Stitch employs a top-down search method to identify new abstractions that can compress the corpus. However, the programs in the corpus are not represented in a normal form. Consequently, subprograms may be semantically equivalent and thus qualify to compress the corpus, e.g., $t_{1}(x)=x+1$ and $t_{2}(x)=1+x$. However, as they are written in different syntactical forms, they are hard to identify.
Programs can be rewritten in a normal form using defined rules. Still, as the number of transformation rules increases and the frequency of their application grows, rewriting the corpus becomes increasingly impractical. This is in parallel to the search space blow-up in program synthesis. For this reason, Stitch does not use a set of predefined rewrite rules, but sacrifices the completeness of compression for a more efficient approach: Stitch essentially employs a program synthesis algorithm that does not aim to solve given tasks but to generate programs that compress the program corpus.
$f_{1}(x)$ $=Add(3,Subtract(Multiply(2,3),1))$
$f_{2}(x)$ $=Add(3,Subtract(Add(1,2),Multiply(3,Subtract(1,4))))$
$a_{par}(x)$ $=Add(3\>,\>??)$
Stitch begins with an empty abstraction (or program) and sequentially adds operations from the DSL. When an operation is added to an abstraction, its arguments are empty or rather contain placeholders. In the next iteration, these placeholders can be filled with new operations, variables, or constants. To make abstractions more general, Stitch allows partial abstractions ($a_{par}$). Partial abstractions introduce new variables to the abstraction - termed ’holes’ (??????? ?) - which can be used to ’fill’ placeholders. Given a program corpus consisting of $f_{1}(x)$ and $f_{2}(x)$ and the partial abstraction $a_{par}(x)$ that was generated in the previous iterations, Stitch iterates over the actions in the DSL and compares the resulting abstractions to the program corpus. In this example, a possible extension would be $a_{par}(x)=Add(3\>,Subtract(\cdot,\cdot))$ which can be further refined in the next iterations.
Iterating over all operations would lead to an exhaustive top-down search which is not feasible. This again relates to the breadth and depth of the search tree. For this reason, Stitch selects actions that compress the program corpus most, i.e., operations that essentially maximize the product of the number of occurrences of an abstraction in the program corpus and the length of the abstraction. If the criterion of an operation added to the current abstraction is below the criterion of the current abstraction itself, the entire branch can be safely pruned. As a consequence of this aggressive pruning strategy, significant gains in processing speed and memory efficiency are achieved.
3 AbstractBeam Framework
To highlight the potential that AbstractBeam has over LambdaBeam, assume a DSL and a target program $p(l)$ that doubles all even elements in a list $l$ and drops all odd ones.
$DSL$ $=[IsEven(x),\>Double(x),\>If(x,then,else),\>Loop(l,start,stop),\>Len(l)]$
$constans$ =[0, 1, 2, 3]absent0123=[0,\>1,\>2,\>3]= [ 0 , 1 , 2 , 3 ]
$p(l)$ $=Loop(l,0,Len(l))(If(IsEven(x),Double(x),Drop(x)))$
Using an exhaustive search, more than $|DSL|^{programlength}=8^{9}>130$ million candidate programs have to be tested to find the solution. If LambdaBeam always suggests the correct operation, a single candidate of weight 8 is generated. As this assumption is likely to be incorrect, assume that LambdaBeam only struggles to suggest the start and end of the loop operation. With this assumption, 5 options, i.e., the constants and $Len(l)$, are available for both start and end yielding $322^{5}=$ 32 candidate programs that have to be executed to find the solution which is more than 7 magnitudes lower than with the primitive approach. As many list manipulations require looping over the entire list, assume that AbstractBeam found an abstraction $a(l)$ in a previous iteration. With the same assumption as for LambdaBeam, AbstractBeam consequently only suggests a single candidate program $p_{AB}(l)$ of length 5 that solves the task.
$a(l)$ $=Loop(l,0,len(l))$
$p_{AB}(l)$ $=a(l)(If(IsEven(x),Double(x),Drop(x)))$
This toy example illustrates the power of LambdaBeam in finding a solution with a few candidate programs only. However, AbstractBeam enhances this property even. Additionally, LambdaBeam is not efficient in terms of time which is a property that is also improved in AbstractBeam as the program length decreases when using abstractions.
AbstractBeam
As shown in Figure 3, AbstractBeam employs a wake-sleep learning cycle similar to the one in DreamCoder [15].
Figure 3: Systematic overview of wake-sleep cylce of the AbstractBeam framework. During the wake phase, the programs are generated to solve given tasks. Solutions to these tasks are then used in the first sleep phase to find abstractions. Finally, the model architecture and the training data is adapted to newfound abstractions and the model is trained to adapt to the new DSL.
In the wake phase, LambdaBeam’s search policy generates programs to solve the training tasks. In order to guarantee enough programs for the Library Learning algorithm, the Unique Randomizer technique and random restarts are used not only during testing, contrary to LambdaBeam, but also during the wake phase. The sleep phase is split in two: First, Stitch identifies possible abstractions from the solutions found in the wake phase. Task solutions are stored across iterations but only the two shortest solutions are given to Stitch to identify potential abstractions. Moreover, abstractions cannot be built based on programs solving the same task and must contain at least two values. This prevents abstractions from being task-specific. Moreover, abstractions like $Add(x2,x1)$ are not added to the DSL avoiding an undesired bloating of the DSL. After refining and adding them to the DSL, the synthesis model is trained in the second phase to learn to use these newfound operations.
Contrary to DreamCoder, the model is not trained on replayed experiences from the wake phase but solely on random tasks constructed from the current DSL. As each cycle’s DSL contains abstractions that were found in the previous iteration, an enhanced DSL produces more complex random tasks which the model then learns from to solve more tasks during the wake phase. These solutions, in turn, help to create a refined DSL in the next iteration. However, this means that the random tasks contain abstractions that were found on the real dataset. For this reason, we need to split the dataset into a training and test set to make sure that no information from the test set is leaked into the training process. In contrast to training LambdaBeam where training samples can be pre-generated by performing exhaustive bottom-up searches without neural guidance once before running and training the synthesizer, for AbstractBeam training data need to be generated at the start of each iteration. This is necessary because new abstractions are possibly added to the DSL which requires adapting the training data to accommodate these new operations. Additionally, the Argument Selector Module of the model must be updated as each operation entails its own module. Consequently, new operation-specific LSTMs must be incorporated for each added operation. All remaining modules can be reused without adaptions. To bootstrap the training of newly added Argument Selector models, they were initialized with the weights of the outermost operation used in the abstraction. Especially the repeated training data generation causes a considerable overhead. Nevertheless, both sleep phases help manage the combinatorial explosion in search: Higher-level DSLs simplify problem-solving with fewer function calls, thus reducing search depth, while the LambdaBeam search policy prioritizes more probable paths, narrowing the search breadth.
To utilize Stitch, programs generated by the bottom-up synthesis need to be parsed into a Lisp-like lambda calculus syntax. Additionally, programs must use the de Brujin Notation [13]. Following this notation programs can be defined as a nested composition of multiple functions where each function is applied to the arguments inside its own parentheses. Variables are replaced by integer indices which refer to the binding depth of lambda expressions, i.e., $\lambda x.(\>\lambda y.(\>x\>+\>y))$ is rewritten to $(lam\>(lam\>(+\>\$1\>\$0)))$. After an abstraction is found by Stitch it is converted back into the original format. As AbstractBeam employs a bottom-up program synthesis, it requires abstractions to be directly executable. However, Stitch is also building partial abstractions that are not directly executable. To account for this, holes in partial abstractions are replaced by variables so when given the variable the operation is executable. These variables count as additional arguments to the operation. Finally, abstractions that do not contain any variables can be executed and are stored as constants.
4 Evaluation
4.1 Experimental setup
Configurations
To manage the volume and complexity of training tasks, we imposed a 1000-second time limit on each data generation episode, during which tasks with a maximum weight of 15 were created. The data generation timeout was increased by 100 seconds per added operation. This clear increase was taken since during training data generation a complete bottom-up search is conducted and consequently, larger DSLs lead to an exponential increase of search duration. Up to 300 searches were conducted in parallel to generate training tasks. In the wake phase, a time limit of 100 seconds per task was used. Random restarts were performed every 10 seconds. Furthermore, both algorithms were trained for a maximum of 10000 steps per iteration. We chose those parameters to be lower than the ones used in [34] to account for fewer available resources. Both approaches were trained for 10 iterations using four RTX 2080 Ti GPUs. The models performing best on the training tasks were used for evaluation. We also train LambdaBeam for 10 iterations to avoid any adverse effects that put LambdaBeam at a disadvantage. 50 handwritten tasks were used as a training set. Note that those 50 tasks were only given to the model during the wake phase to potentially define abstractions and not to train the model. A two-fold cross-validation was performed to cancel out any distribution effects. For each test set, the evaluation was performed five times. Student’s t-tests (significance level 5%) were performed to identify significant differences between both approaches after checking for equal variances and normal distributions. Error bars are displayed as minimum and maximum across trials.
Dataset & DSL
The dataset contains a mixture of 200 tasks — 100 handwritten and 100 synthetic — each designed to be challenging and varied. While the handwritten tasks were designed by hand, the synthetic tasks were created by performing exhaustive bottom-up searches. The synthetic dataset allows us to test whether our hypothesis that abstractions are only helpful if they reflect the domain and consequently whether LambdaBeam’s performance acts as a lower bound to AbstractBeam’s performance. Both datasets are the same that were used to evaluate LambdaBeam [34]. Each task includes between one and three input variables and three to five I/O examples. We used the LambdaBeam DSL but with constants $-1,1$ to assess the boosting of LambdaBeam due to Library Learning. Moreover, we also used the original LambdaBeam DSL to assess the potential of Library Learning in domains not designed for finding abstractions. Exemplary tasks can be found in Appendix A and details on the DSLs in Appendix C.
4.2 Results
LambdaBeam solved $50.2\%\pm 2.0\%$ of all handwritten test tasks when trained on the first fold and $36.0\%\pm 2.5\%$ on the second fold. AbstractBeam significantly outperformed LambdaBeam achieving $53.6\%\pm 2.3\%$ ($p<0.05$) on the first and $39.6\%\pm 1.5\%$ ($p<0.05$) on the second fold. This confirms our first hypothesis that Library Learning leads to an improved performance of LambdaBeam’s synthesis. Exemplary task solutions can be found in Appendix A and a list of all found abstractions in Appendix D.1.
((a)) Success rate per program length
((b)) Abstraction usage per program length
Figure 4: AbstractBeam outperforms LambdaBeam. The difference is especially dominant for program length 7. Here, 62% and of programs contain abstractions. The differences are all significant for program lengths that use abstractions except length 8. ${ns:p>=0.05},{*:p<0.05},{***:p<0.001},{****:p<0.0001}$
As depicted in Figure 4, significant performance differences can only be found for program lengths that use abstractions. For program lengths where no or only a few program solutions contain abstractions, the performance does not differ significantly between both approaches. This also holds true for the second fold (Figure 8). This further confirms that the overall performance boost is caused by the use of abstractions.
We hypothesized that the abstraction-induced search depth reduction causes the performance boost and addresses LambdaBeam’s weakness of low performance in terms of time.
((a)) Enumeration time
((b)) Candidate programs
Figure 5: With less enumeration time and fewer candidate programs AbstractBeam achieves better performance than LambdaBeam.
To further analyze this, Figure 5(a) shows the performance of both approaches over the enumeration time, i.e., the elapsed time until a solution was found. Clearly, AbstractBeam shows an increased performance with less enumeration time. This proves the decrease of the search space depth in AbstractBeam. Moreover, AbstractBeam also shows an improved performance with fewer candidate programs (Figure 5(b)). This also relates to the reduction of the search space depth and boosts LambdaBeam’s proficiency even further. Interestingly, AbstractBeam enumerates fewer programs than LambdaBeam even though both algorithms used the same timeout to generate programs. This relates to the fact that adding abstractions to the DSL lowers the search depth but increases its width. With an increased search width, more candidate programs are suggested. Yet, as AbstractBeam still outperforms LambdaBeam, the balance between a larger DSL and a reduced search depth is advantageous. However, adding abstractions that are not used and bloat the DSL will shift this balance and neglect all benefits of Library Learning. This highlights the care that must be taken when designing the initial and altering the current DSL.
To assess the degree to which abstractions scale beyond the domain they were found in, we evaluated both approaches on the synthetic dataset. No significant performance differences were found between AbstractBeam (fold 1: $52.6\%\pm 1.4\%$, fold 2: $52.4\%\pm 1.9\%$) and LambdaBeam (fold 1: ${53.8\%\pm 1.2\%}$, fold 2: $51.2\%\pm 1.7\%$).
((a)) Success rate per program length
((b)) Abstraction usage per program length
Figure 6: On the synthetic dataset, no significant performance differences between LambdaBeam and AbstractBeam exist. This is because abstractions found on the handwritten dataset do not reduce the search depth in the synthetic domain. $ns:p>=0.05$
Even though abstractions are used in the solution of synthetic tasks, they do not reduce the search depth (Figure 6). This is in accordance to our hypothesis because the synthetic dataset does not reflect the distribution of programs and tasks of the handwritten dataset (Figure 11). Consequently, the abstractions found on the handwritten dataset do not fit into the synthetic domain. Finally, we trained both approaches using the original LambdaBeam DSL and constants. The findings are in parallel to the adapted DSL: AbstractBeam ($44.8\%\pm 2.0\%$) significantly outperformed LambdaBeam ($39.6\%\pm 0.8\%$) . More details can be found in Appendix D.2.
Limitations
The key limitation of AbstractBeam is the structure of the domain which imposes two constraints: First, Library Learning only works if abstractions are possible. This requires that sufficient program repetitions must exist in the solution space and sufficient tasks (or data samples) must contain these program repetitions. Second, on a more practical note, the model must be able to solve sufficient tasks, especially in the first iteration. Otherwise, the Library Learning algorithm cannot identify program repetitions even though sufficient tasks exist that contain a potential abstraction. While both points are restrictions of the task domain, the latter also relates to the design of the initial DSL that must be specific enough to solve tasks but general enough to allow repetitions.
5 Related work
Human behavior has been a fundamental source of inspiration to develop new approaches in Machine Learning ranging from the development of the first ANNs [19] to Generative Agents imitating human beings in social interactions [27]. In program synthesis, this idea is also pervasive. While some approaches train ANNs on solution traces gathered from human beings [25], others mimic human strategies by employing hierarchical approaches that divide problems into smaller, easier tasks [21, 22, 17, 35, 41, 37, 40].
These techniques aim to develop new solution strategies. Yet, they do not devote themselves to tackling the main bottleneck of program synthesis, i.e., the exploding search space. A common approach to reducing the search space is by guiding a search algorithm towards more promising regions. Methods differ in the way of guidance ranging from guidance via prior knowledge [39] to executing intermediate programs to condition the subsequent search on [8, 42, 14, 9, 23]. In general, ANNs are employed to identify auspicious areas [39, 14, 20]. While guiding the search effectively reduces the search breadth, other approaches focus on reducing the search depth by extending the initial DSL with programs that were commonly used in previous programs [15, 18, 11, 24]. This strategy also sparked interest in more efficient algorithms to identify common subprograms [6, 5].
Our approach uses a framework inspired by DreamCoder [15] yet builds upon LambdaBeam’s [34] search strategy. LambdaBeam itself adapts CrossBeam [33] to allow the use of higher-order functions. Both, employ multi-level, neural, and execution-guided synthesis. By integrating LambdaBeam into an architecture inspired by DreamCoder we add Library Learning as a fourth strategy to reduce the search space. Opposed to LambdaBeam, DreamCoder employs a top-down search strategy and does not condition its program generation on intermediate execution information. Moreover, the DreamCoder DSL consists of primitive operations yielding many commonly used subprograms. Furthermore, DreamCoder’s list domain is manually designed to show the effect of DSL enhancement by consisting of task clusters differing in one parameter only, e.g., add-k with k ranging from zero to five. Finally, LambdaBeam does not use any form of Library Learning while DreamCoder uses a different algorithm than Stitch. To the best of our knowledge, this is the first work assessing the effect of DSL enhancement combined with a search algorithm that uses bottom-up execution-guided search.
6 Conclusion
We introduce AbstractBeam, an execution-guided program synthesis framework employing Library Learning. Consequently, AbstractBeam enhances its DSL in a data-driven manner based on the domain it is applied to. Our results show that as a consequence of this an improved performance and a more efficient program generation are achieved. As Library Learning is data-driven, abstractions do not generalize beyond the domain which is also reflected in our experimental results. Consequently, the hurdle of designing a fitting DSL remains. Adding domain-specific abstractions to the DSL likely also incorporates domain knowledge in the DSL which helps the model to understand the domain better. However, the degree to which this support program synthesis is yet to be determined.
Appendices
Appendix A Dataset
Exemplary tasks from the handwritten dataset taken from LambdaBeam [34] are displayed below.
A.1 Task “count:num-under-first”
To solve this task, elements lower than the first element in each input list must be counted.
1inputs_dict = {
2’x’: [[5, 2, 6, 7, 1, 3, 8],
3 [26, 0, 19, 22, 39, 47, 1, 26],
4 [65, 33, 16, 53, 34, 95, 86, 4, 12],
5 [-11, 23, -48, -24, -11, -4, 44, 7, -26, -12],
6 [-11, 23, -48, -24, -11, -14, 44, 7, -26, -12]]
7}
8outputs = [3, 4, 6, 4, 5]
9solution=’Count((lambda u1: Less(u1, Head(x))), x)’
AbstractBeam used this program to solve the task:
1 fn_0(Head(x), x)
2 fn_0 = Count(Less(v1, x1) x2)
Here, the solution has only weight four and is generated in $0.35\pm 0.14$ seconds. LambdaBeam’s solution has weight seven and was only found in two of five runs ($9.86$ and $22.01$ seconds).
1 Count((lambda u1: ((lambda v1: Greater(Head(x), v1)))(u1)), x)
A.2 Task “filter:negative-descending”
This task filters all negative numbers in the input list $x$ and sorts them in descending order.
1inputs_dict = {
2 ’x’: [[-4, 8, 1, -2, -9, -5],
3 [-11, 10, -13, -19, 0, -12, -10],
4 [3, -38, -48, 32, -43, -19, 23, -36, -44, 48]]
5 }
6outputs = [[-2, -4, -5, -9],
7 [-10, -11, -12, -13, -19],
8 [-19, -36, -38, -43, -44, -48]]
9solution = ’Filter((lambda u1: Less(u1, 0)), Reverse(Sort(x))))’
Interestingly, AbstractBeam found an alternative solution in $0.69\pm 0.27$ seconds which is also of weight seven by using an abstraction:
1 Reverse(Take(fn_0(-1, x), Sort(x)))
2 fn_0 = Count(Less(v1, x1) x2)
With this program, the input list is first sorted in ascending order. Then, the number of elements smaller than -1 is determined. Finally, the first n elements are taken from the sorted list and reversed. LambdaBeam also found an alternative solution of weight 8 in $1.58\pm 0.34$ seconds.
1Reverse(Sort(Filter((lambda u1: ((lambda v1: Greater(-1, v1)))(u1)), x)))
Appendix B LambdaBeam
The bottom-up neurally-guided search algorithm, LambdaBeam [34], is an enhancement of its precursor CrossBeam [33]. LambdaBeam was selected for multiple reasons: its bottom-up construction of programs mimics human programming behavior and facilitates the integration of semantic information within the search. Furthermore, it supports the use of higher-order functions in the DSL, thereby expanding its capability to address various tasks. LambdaBeam is recognized for its superior performance as a state-of-the-art, bottom-up neurally-guided search algorithm.
Building higher-order lambda expressions
When generating programs using a bottom-up search approach in lambda calculus, a common problem is the inadvertent creation of semantically equivalent programs, i.e., program aliasing. This redundancy not only bloats the search space but can also lead to inefficiencies in the synthesis. For example, the terms $t_{1}=\lambda x.x+y$ and $t_{2}=\lambda x.y+x$ both perform the same operation but are structurally different. Additionally, $y$ is a free variable in both terms. This leads to ambiguity in the terms’ semantics since $y$ is not bound within the term, and thus its value is undefined and depends on the context in which the term is invoked. Recognizing and reducing such equivalences is both crucial for streamlining the synthesis process and facilitating the assembly of complex lambda functions from simpler components, which aids their integration during the bottom-up search process.
To address the issues of program aliasing and program ambiguity, LambdaBeam employs a specialized approach called the MERGE operator. This operator is designed to construct larger, more complex lambda expressions from simpler, previously created terms while ensuring that all intermediate terms do not contain free variables. The MERGE operator systematically includes additional lambda expressions for each argument that is a function. These supplemental lambdas use variables labeled as $u_{i}$. Any other variables that serve to unify variables across different arguments are labeled $v_{i}$. This procedure aligns with a recognized standard in typed lambda calculus called $\eta$-long normal form [28]. An example of this can be seen in Equation 1 and Equation 2 where $x$, $f$, and $r$ are input variables containing lists.
$t_{before}=\text{Map}\left(\lambda u_{1}:\text{If}\left(\text{Equal}(f,u_{1}),r,u_{1}\right),x\right)$ (1)
$t_{after}=\text{Map}\left(\lambda u_{1}:\left(\lambda v_{1}:\text{If}\left(\left(\lambda v_{1}:\text{Equal}(f,v_{1})\right)(v_{1}),r,v_{1}\right)\right)(u_{1}),x\right)$ (2)
MERGE ensures that each combined term conforms to the DSL’s syntax and semantics without introducing any new free variables. It pads expressions with additional lambda abstractions as needed to bind any new or unbound variables and canonicalizes variable naming and ordering (Equation 2). This process not only ensures syntactic correctness but also maintains semantic integrity by preserving the functional behavior of the combined terms.
Learning
A key challenge in program synthesis is to allow Neural Networks to reason about programs and operations. To address this challenge, property signatures are used in LambdaBeam which assign a set of properties to each expression. Unary properties characterize individual values by associating them with a Boolean condition, for instance, whether a list is empty. Binary properties capture the relationships between inputs and outputs, such as, whether the length of both lists is the same. Property signatures for non-lambda expressions can be created by evaluating the expressions on actual I/O examples. On the contrary, the direct evaluation of lambda expressions is often not feasible, as the specific application context and argument structure in the final solution remain undefined. Therefore, lambda expressions are evaluated on predefined sets of canonical argument tuples, carefully chosen to match the number and types of arguments, ensuring uniform training and evaluation outcomes. The results are captured using unary property signatures and compared with outputs and individual arguments through binary properties, aiding the model in understanding task requirements and argument interactions. Property vectors from multiple runs are condensed to calculate the prevalence and validity of each property, creating an extensive property signature.
Appendix C DSL
We used the same DSL as in LambdaBeam [34]. This consists of the constants $-1,0,1,2,3,4$, 23 first-order, and five higher-order functions. Our adapted DSL just contained the constants $-1$ and $1$. Other than that both DSLs are identical.
# 23 first-order operations
def Add(x: int, y: int)-> int:
return x + y
def Subtract(x: int, y: int)-> int:
return x - y
def Multiply(x: int, y: int)-> int:
return x * y
def IntDivide(x: int, y: int)-> int:
return x // y
def Square(x: int)-> int:
return x ** 2
def Min(x: int, y: int)-> int:
return min(x, y)
def Max(x: int, y: int)-> int:
return max(x, y)
def Greater(x: int, y: int)-> bool:
return x > y
def Less(x: int, y: int)-> bool:
return x < y
def Equal(x: int, y: int)-> bool:
return x == y
def IsEven(x: int)-> bool:
return x % 2 == 0
def IsOdd(x: int)-> bool:
return x % 2 != 0
def If(c: bool, x: int, y: int)-> int:
return x if c else y
def Head(xs: list)-> int:
return xs[0]
def Last(xs: list)-> int:
return xs[-1]
def Take(n: int, xs: list)-> list:
return xs[:n]
def Drop(n: int, xs: list)-> list:
return xs[n:]
def Access(n: int, xs: list)-> int:
return xs[n]
def Minimum(xs: list)-> int:
return min(xs)
def Maximum(xs: list)-> int:
return max(xs)
def Reverse(xs: list)-> list:
return list(reversed(xs))
def Sort(xs: list)-> list:
return sorted(xs)
def Sum(xs: list)-> int:
return sum(xs)
# 5 higher-order operations
def Map(f: Callable[[int], int], xs: list)-> list:
return [f(x) for x in xs]
def Filter(f: Callable[[int], bool], xs: list)-> list:
return [x for x in xs if f(x)]
def Count(f: Callable[[int], bool], xs: list)-> int:
return len([x for x in xs if f(x)])
def ZipWith(f: Callable[[int, int], int], xs: list, ys: list)-> list:
return [f(x, y) for x, y in zip(xs, ys)]
def Scanl1(f: Callable[[int, int], int], xs: list)-> list:
ys = [xs[0]]
for n in range(1, len(xs)):
ys.append(f(ys[n-1], xs[n]))
return ys
Appendix D Supplemental Results
D.1 Abstractions
AbstractBeam found 8 abstractions when trained on the first fold and 11 abstractions when trained on the second fold. Noteworthy, abstractions do not only contain primitive DSL functions, such as $fn_{3}$ in the first fold, but also lambda functions, e.g., $fn_{0}$ in the first fold. Moreover, abstractions of previously defined abstractions, for instance, $fn_{10}$ in the second fold, and higher-order abstractions, i.e., abstractions that take a function as an input argument, e.g., $fn_{1}$ in the first fold, were found. All found abstractions are displayed in Table 1.
Table 1: Abstractions found on by AbstractBeam. Interestingly, the adaption of the constants to just $-1,1$ did not cause AbstractBeam to abstract the constants (only in the first fold AbstractBeam abstracted 00).
Fold Abstractions
1 $constant:0$
$fn_{0}=Count((lambda\>v1:Less(v1,x1)),x2)$
$fn_{1}=x1(Min(Add(1,1),x2))$
$fn_{3}=Add(-1,x1)$
$fn_{5}=Count((lambda\>v1:Equal(v1,x1)),x2)$
$fn_{6}=x1(Min(fn_{3}(1),x2))$
$fn_{8}=x1(Min(0,x2))$
$fn_{14}=x1(0,x2)$
2 $fn_{0}=Reverse(Sort(x1))$
$fn_{1}=Take(-1,x1)$
$fn_{3}=Sum(Take(x2,x1))$
$fn_{9}=ZipWith((lambda\>u1,\>v1:(Subtract(v1,x1))),x3,x2)$
$fn_{10}=fn_{9}(Minimum(x1),x1,x1)$
$fn_{11}=Sum(ZipWith((lambda\>u1,\>v1:(Multiply(v1,u1))),x2,x1))$
$fn_{14}=Scanl1((lambda\>u1,\>v1:(Add(v1,u1))),x1)$
$fn_{18}=ZipWith((lambda\>u1,\>v1:(Multiply(v1,u1))),x2,x1)$
$fn_{21}=Filter((lambda\>v1:IsOdd(v1)),x1)$
$fn_{22}=Scanl1((lambda\>u1,\>v1:(Multiply(v1,u1))),x1)$
$fn_{24}=Head(x1(fn_{0}(x2)))$
However, as shown in Figure 7, not all abstractions were used in the generation of solutions.
((a)) Fold 1
((b)) Fold 2
Figure 7: Histogram of the occurrences of DSL operations including abstractions in solutions.
This has several potential reasons. Most likely, the abstractions were not present in the solutions of the test tasks.
D.2 Further Analysis
The findings of the first fold also hold true for the second fold: AbstractBeam significantly outperforms LambdaBeam. As shown in Figure 8, this confirms for the performance in terms of time and in terms of enumerated programs as well.
((a)) Enumeration time
((b)) Candidate programs
((c)) Performance per program length
((d)) Abstraction usage per program length
Figure 8: Performance analysis of AbstractBeam and LambdaBeam on the second fold of the handwritten dataset.
Again, AbstractBeam shows clear performance increases with lower enumeration times and fewer explored values. Furthermore, LambdaBeam enumerates more values in total. This is in parallel to the findings on the first fold and is likely caused by the narrower search tree. Even though, the overall performance of AbstractBeam shows significant differences to LambdaBeam’s performance, the findings are not as clear as in the first fold when looking at the performance per program length. Here, AbstractBeam shows improved performances for most program lengths. However, the performance differences are not significant for all lengths except 11 and 12. AbstractBeam does not use any abstractions in programs of length 11 which explains the better performance of LambdaBeam. Operations and how frequently they were used in solutions are depicted in Figure 9.
((a)) Fold 1
((b)) Fold 2
Figure 9: Histogram of operations used in the solutions.
Similarly to the evaluation using the original DSL and adapted constants $-1,1$, we trained and evaluated AbstractBeam and LambdaBeam using the original DSL and original constants $-1,0,1,2,3,4$. AbstractBeam significantly outperformed LambdaBeam. The performance metrics are shown in Figure 10. Noteworthy, LambdaBeam performs worse than the results stated in [15]. The reason for this is the lower enumeration timeout we used (100s vs. 600s) and fewer resources used to train and evaluate both approaches. However, in parallel to the findings using the adapted constants, AbstractBeam again shows improved performance with less enumeration timeout and fewer candidate programs.
((a)) Abstraction usage per program length
((b)) Performance per program length
Figure 10: Performance analysis of AbstractBeam and LambdaBeam on the synthetic dataset.
No significant performance differences were found when evaluating both approaches on the synthetic dataset. The distributions of the synthetic and the handwritten dataset differ clearly as shown in Figure 11. This explains the performance of AbstractBeam on the synthetic dataset. Here, the abstractions found on the handwritten training set do not occur in the solution space. Thus, they cannot be usefully applied.
((a)) Fold 1
((b)) Fold 2
Figure 11: Histogram of operations in the handwritten and synthetic dataset.
((a)) Enumeration time
((b)) Candidate programs
((c)) Performance per program length
((d)) Abstraction usage per program length
Figure 12: Performance analysis of AbstractBeam and LambdaBeam on the first fold of the synthetic dataset.
As abstractions do not fit the distribution of the synthetic dataset, the performance and efficiency boost that Library Learning gives in AbstractBeam is eliminated. This does not mean that abstractions are not applied in the solution of synthetic tasks but rather that abstractions do not reduce the program length. Consequently, the search depth is not decreased. This is depicted in Figure 12 and Figure 13.
((a)) Enumeration time
((b)) Candidate programs
((c)) Performance per program length
((d)) Abstraction usage per program length
Figure 13: Performance analysis of AbstractBeam and LambdaBeam on the second fold of the synthetic dataset.